# A Tutte-type canonical decomposition of 3- and 4-connected graphs

### v2

@@ 14. Appendix A: Circular saws
For each vertex $x\in X$ we construct two xxYY paths $P^{+}_{x},P^{-}_{x}$ inductively, as follows. See Figure˜14.1 for an example.
@@
Figure 14.1. Paths constructed in the proof of Lemma˜14.1. The edges $aa^{\prime}$ and $bb^{\prime}$ are coloured blue, the paths are coloured red.
@@
We show first that $G$ is $k$-connected, and assume for a contradiction that some two vertices $a,b$ of $G$ are separated by a set $S$ of k1k-1 vertices. Let $a^{\prime}$ be a neighbour of $a$ outside $S$, and let $b^{\prime}$ be a neighbour of $b$ outside $S$. Note that $2(k-1)\geqslant k+1$ since $k\geqslant 3$. Hence by Lemma˜14.1, there are $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $|S|\leqslant k-1$. Hence $G$ is $k$-connected.
Next, we assume for a contradiction that $G$ has a $k$-separation $(A,B)$ such that $S(A,B)$ fails to be the neighbourhood of some vertex of $G$. Let $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. Then there are vertices $a^{\prime}\in N(a)\smallsetminus S(A,B)$ and $b^{\prime}\in N(b)\smallsetminus S(A,B)$. Now Lemma˜14.1 yields $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $(A,B)$ has order $k$. ∎
@@ 15. Appendix B: Comparison of strict tri-separations and tri-separations
The strict tri-separation-decomposition (˜2) is different from the tri-separation decomposition [CK23, Theorem 1] as can be seen in Figures˜15.1 and 15.2. However, the strict tri-separation-decomposition can be viewed as a refinement of the tri-separation decomposition via Lemma˜15.1.

### v3

@@ 13. Appendix A: Circular saws
For each vertex $x\in X$ we construct two xxYY paths $P^{+}_{x},P^{-}_{x}$ inductively, as follows. See fig.˜13.1 for an example.
@@
Figure 13.1. Paths constructed in the proof of lemma˜13.1. The edges $aa^{\prime}$ and $bb^{\prime}$ are coloured blue, the paths are coloured red.
@@
We show first that $G$ is $k$-connected, and assume for a contradiction that some two vertices $a,b$ of $G$ are separated by a set $S$ of k1k-1 vertices. Let $a^{\prime}$ be a neighbour of $a$ outside $S$, and let $b^{\prime}$ be a neighbour of $b$ outside $S$. Note that $2(k-1)\geqslant k+1$ since $k\geqslant 3$. Hence by lemma˜13.1, there are $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $|S|\leqslant k-1$. Hence $G$ is $k$-connected.
Next, we assume for a contradiction that $G$ has a $k$-separation $(A,B)$ such that $S(A,B)$ fails to be the neighbourhood of some vertex of $G$. Let $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. Then there are vertices $a^{\prime}\in N(a)\smallsetminus S(A,B)$ and $b^{\prime}\in N(b)\smallsetminus S(A,B)$. Now lemma˜13.1 yields $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $(A,B)$ has order $k$. ∎
@@ 14. Appendix B: Comparison of strict tri-separations and tri-separations
The strict tri-separation-decomposition (˜2) is different from the tri-separation decomposition [CK23, Theorem 1] as can be seen in figs.˜14.1 and 14.2. However, the strict tri-separation-decomposition can be viewed as a refinement of the tri-separation decomposition via lemma˜14.1.
@@
Acknowledgement.
We thank a referee for valuable comments on the introduction. We thank another referee for bringing [LPPS22] to our attention. We thank Joseph Devine for exciting discussions. We thank Agelos Georgakopoulos for telling us about a theorem of Godsil and Royle that we use in [KP26].